Multiply using Foil Method * 4) (5v - 5)(3v- 6) A) 15v2 + 30 B) 15v2 - 45v + 30 C) 15v2 - 15v- 30 D) 15v2 + 15v- 30 Uu
Answer #1

4 (Sv.5) (31-6) firsts: Sv (3r) = 15v2 outsides : 50X(-6)=-30v Insides : -5(30) = 150 Lasts: -50 x-6=+3000 ** (5V-5) (3v-6) =